# On the geometry of punctual Hilbert schemes on singular curves and their motivic zeta functions

### v3

@@ 4. Piecewise fibrations induced by the edges of the $\Gamma$-subsemimodules tree
2.We define the syzygy set of $\Delta$ as
@@
[14] We consider $\Gamma=\langle p,q\rangle$ with $p<q$ coprime. Let $\Delta$ be a subsemimodule of $\Gamma$. There exists a bijection between $ASyz(\Delta)$ and $T_{Syz(\Delta)}$ — the minimal generating system of $Syz(\Delta)$. Specifically, the projection
@@
By the Chinese remainder theorem (or see also Section 5 of [15]), every element $m\in\Gamma$ admits a unique presentation of the form $m=aq+bp$ with $0\leq a<p$.
@@
Proposition 4.2 allows us to equate $ASyz(\Delta)$ with $T_{Syz(\Delta)}$ for the semigroup $\Gamma=\langle p,q\rangle$ in the rest of the paper.
Now we will study the canonical morphism between $C^{[\Delta]}$ and $C^{[m(\Delta)]}$ given by an edge, $\Delta\rightarrow m(\Delta)$, in the tree $G_{\Gamma}$.
Let $(C,O)$ be the germ of a unibranch curve singularity with value semigroup $\Gamma$, where $\Gamma$ is either monomial in the sense of [16] (see Section 6) or of the form $\langle p,q\rangle$ with $\gcd(p,q)=1$. We write $\Gamma=\langle\alpha_{1},\dots,\alpha_{e}\rangle$. Let $\Delta=(\gamma_{1},\dots,\gamma_{n})_{\Gamma}$ be a $\Gamma$-subsemimodule, and denote by $\operatorname{Syz}(\Delta)=(\sigma_{1},\dots,\sigma_{m})_{\Gamma}$ its syzygy subsemimodule. Let
@@ 4.1. Proof of the main theorem
By the proof of [14, Theorem 13], the linear forms $L_{i}^{r}$ (and similarly $(L_{i}^{r})^{\prime}$) are linearly independent. Since the higher-order terms in $Eq_{i}^{r}$ are in the ideal generated by the $\lambda,\nu$ variables, the zero locus of $Eq_{i}^{r}=0$ is analytically isomorphic to the zero locus of $L_{i}^{r}=0$. Hence, the strata are defined by their linear parts.
@@
As an application of Theorem 4.1, we recover [14, Theorem 13]:
@@
$C^{[\Delta]}\cong\operatorname{Spec}\mathbb{C}[\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{1}^{10},\lambda_{2}^{7},\lambda_{1}^{23},\lambda_{2}^{20},\lambda_{3}^{10},\lambda_{4}^{7},\lambda_{1}^{29},\lambda_{2}^{26},\lambda_{3}^{16},\lambda_{4}^{13}]/<\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{4}^{7},\lambda_{2}^{7},\lambda_{2}^{20}+(\lambda_{1}^{10})^{2},\lambda_{1}^{10}-\lambda_{3}^{10}>$
@@
$C^{[m(\Delta)]}\cong\operatorname{Spec}\mathbb{C}[\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{1}^{10},\lambda_{2}^{7},\lambda_{1}^{23},\lambda_{2}^{20},\lambda_{3}^{10},\lambda_{4}^{7}]/<\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{4}^{7}\cdot\lambda_{2}^{7}>$
@@
$C^{[m^{2}(\Delta)]}\cong\operatorname{Spec}\mathbb{C}[\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{1}^{10},\lambda_{2}^{7}]/<\lambda_{2}^{1}>$
@@ 4.2. The motivic Hilbert zeta functions of some plane curve singularities
The Grothendieck ring of complex varieties, denoted $K_{0}(\mathrm{Var}_{\mathbb{C}})$, is the ring generated by isomorphism classes $[X]$ of complex algebraic varieties $X$, subject to the following relations:
  • $[X]=[Y]$ if $X\cong Y$,
  • $X]=[X\setminus Y]+[Y]$ for any closed subvariety $Y\subseteq X$, with multiplication defined by $X]\cdot[Y]=[X\times Y]$
@@ 4.2.1. Case of singularities of types $E_{6}$, $E_{8}$, $W_{8}$, $Z_{10}$
We recall a result from the proof of Lemma 17 in [15]. In the case $\Gamma=\langle k,n\rangle$ with $\gcd(k,n)=1$, the minimal generators of a $\Gamma$-subsemimodule can be explicitly described in terms of the minimal generators of $\Gamma$ itself.
@@
[15] There is a one-to-one correspondence between monomial ideals in $\mathbb{C}[[t^{k},t^{n}]]$ and sequences
@@
[15] Let $1\leq\ell\leq 2\delta$. Then every $\Delta\in\mathscr{D}_{\ell}$ has the form
@@ 5. Subvarieties with fixed minimal number of generators in the punctual Hilbert scheme
In this section, motivated by conjectures of Oblomkov, Rasmussen, and Shende [15, 14], we study the geometry and motivic class of subvarieties of the punctual Hilbert scheme that parametrize ideals of $\mathcal{O}_{C}$ with a fixed minimal number of generators. These conjectures propose deep connections between the geometry of the Hilbert scheme of a plane curve singularity and knot invariants—specifically, the HOMFLY polynomial and Khovanov–Rozansky HOMFLY homology. The first of these conjectures was proved by Maulik in [13].
We begin by recalling the conjectures from [15, 14].
@@
Oblomkov and Shende conjectured the following relationship between the Hilbert scheme and the HOMFLY polynomial of the associated link [15], later proven by Maulik [13]:
@@
In [15, Section 5], let $C$ be the curve defined by $y^{k}=x^{n}$. The authors consider a $\mathbb{C}^{*}$-action on $C^{[\ell],m}$. By [2, Corollary 2], only the fixed points contribute to the Euler characteristic, so
@@
For the curve $C$ defined by $y^{k}=x^{n}$, we have
@@ 6. On Monomial semigroups
We recall the notion of monomial semigroups as introduced in [16].
Definition 6.1 ([16]).
@@
We have the following lemma, which is implicit in the proof of [16, Theorem 11]:
@@
Acknowledgments. The authors are grateful to Eugene Gorsky for discussions related to this work. We extend our deep gratitude to Laurent Evain and Evelia Rosa García Barroso for their meticulous review of the manuscript; their insightful comments were invaluable in refining the final version. The third author also acknowledges financial support from the Guangzhou Elites Sponsorship Council (China) through the Overseas Study Program of the Guangzhou Elite Project.

### v4

@@ 1. We define the Augmented Syzygy of $\Delta$ as
2. We define the syzygy set of $\Delta$ as
@@
[undefm] We consider $\Gamma=\langle p,q\rangle$ with $p<q$ coprime. Let $\Delta$ be a subsemimodule of $\Gamma$. There exists a bijection between $ASyz(\Delta)$ and $T_{Syz(\Delta)}$ — the minimal generating system of $Syz(\Delta)$. Specifically, the projection
@@
By the Chinese remainder theorem (or see also Section 5 of [undefn]), every element $m\in\Gamma$ admits a unique presentation of the form $m=aq+bp$ with $0\leq a<p$.
@@
Proposition 4.2 allows us to equate $ASyz(\Delta)$ with $T_{Syz(\Delta)}$ for the semigroup $\Gamma=\langle p,q\rangle$ in the rest of the paper. Now we will study the canonical morphism between $C^{[\Delta]}$ and $C^{[m(\Delta)]}$ given by an edge, $\Delta\rightarrow m(\Delta)$, in the tree $G_{\Gamma}$.
Let $(C,O)$ be the germ of a unibranch curve singularity with value semigroup $\Gamma$, where $\Gamma$ is either monomial in the sense of [undefo] (see Section 6) or of the form $\langle p,q\rangle$ with $\gcd(p,q)=1$. We write $\Gamma=\langle\alpha_{1},\dots,\alpha_{e}\rangle$. Let $\Delta=(\gamma_{1},\dots,\gamma_{n})_{\Gamma}$ be a $\Gamma$-subsemimodule, and denote by $\operatorname{Syz}(\Delta)=(\sigma_{1},\dots,\sigma_{m})_{\Gamma}$ its syzygy subsemimodule. Let
@@ 4.1. Proof of the main theorem
By the proof of [undefm, Theorem 13], the linear forms $L_{i}^{r}$ (and similarly $(L_{i}^{r})^{\prime}$) are linearly independent. Since the higher-order terms in $Eq_{i}^{r}$ are in the ideal generated by the $\lambda,\nu$ variables, the zero locus of $Eq_{i}^{r}=0$ is analytically isomorphic to the zero locus of $L_{i}^{r}=0$. Hence, the strata are defined by their linear parts.
@@
As an application of Theorem 4.1, we recover [undefm, Theorem 13]:
@@
$C^{[\Delta]}\cong\operatorname{Spec}\mathbb{C}[\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{1}^{10},\lambda_{2}^{7},\lambda_{1}^{23},\lambda_{2}^{20},\lambda_{3}^{10},\lambda_{4}^{7},\lambda_{1}^{29},\lambda_{2}^{26},\lambda_{3}^{16},\lambda_{4}^{13}]/\langle\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{4}^{7},\lambda_{2}^{7},\lambda_{2}^{20}+(\lambda_{1}^{10})^{2},\lambda_{1}^{10}-\lambda_{3}^{10}\rangle$
@@
$C^{[m(\Delta)]}\cong\operatorname{Spec}\mathbb{C}[\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{1}^{10},\lambda_{2}^{7},\lambda_{1}^{23},\lambda_{2}^{20},\lambda_{3}^{10},\lambda_{4}^{7}]/\langle\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{4}^{7}\cdot\lambda_{2}^{7}\rangle$
@@
$C^{[m^{2}(\Delta)]}\cong\operatorname{Spec}\mathbb{C}[\lambda_{1}^{4},\lambda_{2}^{1},\lambda_{1}^{10},\lambda_{2}^{7}]/\langle\lambda_{2}^{1}\rangle$
@@ 4.2. The motivic Hilbert zeta functions of some plane curve singularities
he Grothendieck ring of complex varieties, denoted $K_{0}(\mathrm{Var}_{\mathbb{C}})$, is the ring generated by isomorphism classes $[X]$ of complex algebraic varieties $X$, subject to the scissor relation (also called the cut-and-paste relation):
$X]=[X\setminus Y]+[Y]$
for every closed subvariety $Y\subseteq X$. Multiplication is defined by $X]\cdot[Y]=[X\times Y]$.
@@ 4.2.1. Case of singularities of types $E_{6}$, $E_{8}$, $W_{8}$, $Z_{10}$
We recall a result from the proof of Lemma 17 in [undefn]. In the case $\Gamma=\langle k,n\rangle$ with $\gcd(k,n)=1$, the minimal generators of a $\Gamma$-subsemimodule can be explicitly described in terms of the minimal generators of $\Gamma$ itself.
@@
[undefn] There is a one-to-one correspondence between monomial ideals in $\mathbb{C}[[t^{k},t^{n}]]$ and sequences
@@
[undefn] Let $1\leq\ell\leq 2\delta$. Then every $\Delta\in\mathscr{D}_{\ell}$ has the form
@@ 5. Subvarieties with fixed minimal number of generators in the punctual Hilbert scheme
In this section, motivated by conjectures of Oblomkov, Rasmussen, and Shende [undefn, undefm], we study the geometry and motivic class of subvarieties of the punctual Hilbert scheme that parametrize ideals of $\mathcal{O}_{C}$ with a fixed minimal number of generators. These conjectures propose deep connections between the geometry of the Hilbert scheme of a plane curve singularity and knot invariants—specifically, the HOMFLY polynomial and Khovanov–Rozansky HOMFLY homology. The first of these conjectures was proved by Maulik in [undefl].
We begin by recalling the conjectures from [undefn, undefm].
@@
Oblomkov and Shende conjectured the following relationship between the Hilbert scheme and the HOMFLY polynomial of the associated link [undefn], later proven by Maulik [undefl]:
@@
In [undefn, Section 5], let $C$ be the curve defined by $y^{k}=x^{n}$. The authors consider a $\mathbb{C}^{*}$-action on $C^{[\ell],m}$. By [undefa, Corollary 2], only the fixed points contribute to the Euler characteristic, so
@@
[undefn] For the curve $C$ defined by $y^{k}=x^{n}$, we have
@@ 6. On Monomial semigroups
We recall the notion of monomial semigroups as introduced in [undefo].
Definition 6.1 ([undefo]).
@@
We have the following lemma, which is implicit in the proof of [undefo, Theorem 11]:
@@
Acknowledgments. The authors are grateful to Eugene Gorsky for discussions related to this work. We extend our deep gratitude to Laurent Evain and Evelia Rosa García Barroso for their meticulous review of the manuscript; their insightful comments were invaluable in refining the final version. The second author was partially supported by he ANR-SINTROP. The third author also acknowledges financial support from the Guangzhou Elites Sponsorship Council (China) through the Overseas Study Program of the Guangzhou Elite Project.